# Semi-Bandit Learning for Monotone Stochastic Optimization

### v1

@@ Abstract
Stochastic optimization is a widely used approach for optimization under uncertainty, where uncertain input parameters are modeled by random variables. Exact or approximation algorithms have been obtained for several fundamental problems in this area. However, a significant limitation of this approach is that it requires full knowledge of the underlying probability distributions. Can we still get good (approximation) algorithms if these distributions are unknown, and the algorithm needs to learn them through repeated interactions? In this paper, we resolve this question for a large class of “monotone” stochastic problems, by providing a generic online learning algorithm with $\sqrt{T\log T}$ regret relative to the best approximation algorithm (under known distributions). Importantly, our online algorithm works in a semi-bandit setting, where in each period, the algorithm only observes samples from the r.v.s that were actually probed. Our framework applies to several fundamental problems in stochastic optimization such as prophet inequality, Pandora’s box, stochastic knapsack, stochastic matchings and stochastic submodular optimization.
@@ 1 Introduction
Stochastic optimization problems have been a subject of intense investigation, as they offer a powerful lens through which we can handle uncertain inputs. In stochastic problems, uncertain input parameters are modeled by random variables, which are usually independent. Solutions (or policies) to a stochastic problem are sequential decision processes, where the next decision depends on all previously observed information. This approach has been applied to many domains: optimal stopping (Pandora’s box [Wei79, Sin18] and prophet inequality [SC84, KW12, FSZ21]), stochastic submodular optimization [AN16, GK17, INZ16], stochastic probing [GN13, ASW16, GNS17] and various other stochastic combinatorial optimization problems (e.g., knapsack [DHK16, DGV08] and matching [${}^{+}$, ${}^{+}$12]). A fundamental assumption in all these results is that the underlying probability distributions are known to the algorithm. While the known-distributions assumption holds in some settings, it is not satisfied in many practical applications, e.g., in the absence of historical data. Our main goal in this paper is to relax this assumption and handle stochastic problems with unknown distributions. We start by providing some concrete examples of the problems we consider.
Series testing.
@@
There is a system with $n$ components, where each component $i$ is “working” independently with some probability $p_{i}$. All $n$ components must be working for the system to be functional. Moreover, it costs $c_{i}$ to test component $i$ and determine whether/not it is working. The goal is to test components sequentially to determine whether/not the system is functional, at the minimum expected cost. Clearly, testing will continue until some component is found to be not working, or we have tested all components and found them to be working. In the standard setting, where the probabilities $\{p_{i}\}_{i=1}^{n}$ are known upfront, the greedy policy that tests components in decreasing order of $\frac{1-p_{i}}{c_{i}}$ is optimal [But72].
Pandora’s box.
@@
There are $n$ items, where each item $i$ has an independent random value $\mathsf{X}_{i}$ with a known distribution $\mathcal{D}_{i}$. The realized value of $\mathsf{X}_{i}$ can be revealed by inspecting item $i$, which incurs cost $c_{i}$. The goal is to inspect a subset $S$ of the items (possibly adaptively) to maximize the expected utility $\mathbb{E}\left[\max_{i\in S}\mathsf{X}_{i}-\sum_{i\in S}c_{i}\right]$. We emphasize that a solution to this problem can be quite complex: the choice of the next item to inspect may depend on the realizations of all previous items. Nevertheless, there is an elegant optimal solution to this problem when the distributions $\{\mathcal{D}_{i}\}_{i=1}^{n}$ are known [Wei79].
Stochastic covering knapsack.
There are $n$ items, where each item $i$ has a (deterministic) cost $c_{i}$ and independent random reward $R_{i}$ with known distribution $\mathcal{D}_{i}$. The realized value of $R_{i}$ is only known when item $i$ is selected. Given a target $Q$, the goal is to select items sequentially (and adaptively) until the total reward is at least $Q$. The objective is to minimize the expected cost of selected items. Although the deterministic knapsack problem is already NP-hard, there is an elegant $3$-approximation algorithm for stochastic covering knapsack [DHK16] with known distributions.
@@
In each of the above problems, what if the underlying probabilities/distributions are unknown? Is it still possible to obtain good performance guarantees? In order to address this question formally, we utilize the framework of online learning. Here, the algorithm interacts with an unknown-but-fixed input distribution $\mathcal{D}$ over multiple time periods. [^1] In the single period setting, it is easy to construct examples that rule out any reasonable performance bound for these problems with unknown distributions. In each period $t=1,\cdots T$, the online algorithm comes up with a solution/policy $\sigma^{t}$ to the stochastic problem and receives some feedback based on the performance of $\sigma^{t}$ on the (unknown) distribution $\mathcal{D}$.
@@
The type of feedback received is a crucial component in this learning-based framework. The simplest setting is full feedback, where the algorithm receives one sample from every random variable (r.v.). However, this is unrealistic for stochastic problems because the policy $\sigma^{t}$ in period $t$ only observes the r.v.s corresponding to some subset of the items. For example, in series testing, $\sigma^{t}$ tests the $n$ components in some order until the first non-working component is found: if the first $k-1$ components are working and the $k^{th}$ component is not working then we would only observe the outcomes/samples of the first $k$ r.v.s (and the remaining $n-k$ r.v.s are unobserved). In this paper we consider the more natural setting of semi-bandit feedback, where in each period $t$, the algorithm only receives samples from the r.v.s that the policy $\sigma^{t}$ actually observed.
@@
Our goal is to minimize the expected regret of the online algorithm, which is the difference between the total $T$-period objective of our algorithm and the optimum (which knows the distribution $\mathcal{D}$). Obtaining $o(T)$ regret dependence with respect to $T$ means that our online algorithm, which doesn’t know $\mathcal{D}$, asymptotically approaches the optimum. An additional difficulty arises from the fact that many problems that we consider are NP-hard (e.g., knapsack and submodular optimization). In such cases, we cannot hope to approach the optimum (via a polynomial algorithm). Here, we use the notion of $\alpha$-regret, where the online algorithm approaches the $\alpha$ approximately optimal value. In light of this, we ask: can we transform any offline (approximation) algorithm for a stochastic problem into an online learning algorithm that has low-regret with respect to this offline algorithm?
Our main result is an affirmative answer to this question. We give a general method for transforming any offline $\alpha$-approximation algorithm (with known distributions) to an online learning algorithm with $\alpha$-regret of $\widetilde{O}(\sqrt{T})$. It is well-known that the $\sqrt{T}$ regret bound cannot be improved, even in very special cases. Our method works for a wide range of stochastic problems that satisfy a natural monotonicity condition. This includes several fundamental problems in stochastic optimization such as prophet inequality, Pandora’s box, stochastic knapsack, stochastic matching, stochastic submodular maximization and submodular cover.
@@ 1.1 Problem Setup
Stochastic Optimization.
Consider a stochastic problem $\mathcal{P}$ where the input consists of $n$ items with an independent random variable (r.v.) $\mathsf{X}_{i}$ associated with each item $i\in[n]=\{1,2,\cdots n\}$. There is a known probability distribution $\mathcal{D}_{i}$ for $\mathsf{X}_{i}$; that is, $\mathsf{X}_{i}\sim\mathcal{D}_{i}$. We assume that each $\mathcal{D}_{i}$ has finite support, and denote the set of outcomes of all r.v.s by $O$. (We relax the discrete-distribution assumption in §2.6.) In order to determine the realization of any r.v. $\mathsf{X}_{i}$, we need to probe item $i$.
@@
The cost of policy $\sigma$ depends on the realization $\mathbf{x}$, and we use $f(\sigma,\mathbf{x})\geq 0$ to denote the cost of policy $\sigma$ under realization $\mathbf{x}$. Specifically, $f(\sigma,\mathbf{x})$ depends only on the sequence of probed items $S(\sigma,\mathbf{x})$ and their realizations. We assume that cost is accrued only at terminal/leaf nodes: this is without loss of generality as every policy execution ends at a leaf node. [^2] In some problems, it is natural to accrue cost at each state that the policy goes through. We can easily convert such a cost structure to our setting by placing all the cost accrued from a policy execution at its terminal state. Note that we do not make any assumption on the cost function $f$: it does not have to be linear, submodular etc. Our goal is to find a policy $\sigma$ that minimizes the expected cost $f(\sigma):=\mathbb{E}_{\mathbf{x}}[f(\sigma,\mathbf{x})]$. We also allow for constraints in problem $\mathcal{P}$, which must be satisfied under all realizations (i.e., with probability $1$). Let $\mathcal{C}$ denote the set of feasible policies (that satisfy the constraints in $\mathcal{P}$). Sometimes, we work with randomized policies, where each node in the decision tree corresponds to a probability distribution over items (rather than a single item). Randomized policies are not any stronger than deterministic ones: there is always an optimal deterministic policy for this class of problems. Succinctly, the stochastic problem $\mathcal{P}$ is as follows (maximization problems are defined similarly).
@@
Observe that the number of nodes in policy $\sigma$ may be exponential. So, we are interested in “efficient” policies that can be implemented in polynomial-time for any realization $\mathbf{x}$. For some stochastic problems like series systems and Pandora’s box, efficient optimal policies are known. On the other hand, there are many problems like stochastic knapsack, matching and and set-cover, where optimal policies may not be efficient: in these cases, we will focus on (efficient) approximately optimal policies. We use $\sigma^{*}$ to denote an optimal policy, and denote its expected cost by $\mathtt{OPT}$. An $\alpha$-approximation algorithm for a stochastic problem $\mathcal{P}$ takes as input the probability distributions $\{\mathcal{D}_{i}\}_{i=1}^{n}$ (along with any objective/constraint parameters) and returns a policy that has expected cost at most $\alpha\cdot\mathtt{OPT}$. We use the convention that for minimization problems, the approximation ratio $\alpha\geq 1$, whereas for maximization problems $\alpha\leq 1$.
The Online Semi-Bandit Setting.
In this setting, the distributions $\mathcal{D}_{1},\ldots,\mathcal{D}_{n}$ are unknown. Moreover, we have to repeatedly solve the stochastic problem $\mathcal{P}$ over $T$ time periods. We assume throughout the paper that $T\geq n$. The distributions $\{\mathcal{D}_{i}\}_{i=1}^{n}$ remain the same across all $T$ periods. The goal is to simultaneously learn the distributions and converge to a good policy over time. At time $t\leq T$, we use all prior observations to present policy $\sigma^{t}$. The policies $\sigma^{t}$ may be different for each time $t\in[T]$. We measure our learning algorithm in terms of expected total regret. For minimization problems, we define it as follows (the definition for maximization problems is analogous).
@@
$R(T)=\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\mathbb{E}_{\mathbf{x}^{t}}\left[f(\sigma^{t},\mathbf{x}^{t})-f(\sigma^{*},\mathbf{x}^{t})\right]\right]=\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\mathbb{E}_{\mathbf{x}^{t}}\left[f(\sigma^{t},\mathbf{x}^{t})\right]-T\cdot\mathtt{OPT}.$ (2)
@@
An important component in this learning setup is what type of feedback is received by the algorithm in each round. In the full feedback setting, the algorithm gets to know the entire sample $\mathbf{x}^{t}=(x_{1}^{t},\ldots,x_{n}^{t})$, which could be used to update beliefs regarding all the distributions $\mathcal{D}_{1},\ldots,\mathcal{D}_{n}$. However, as mentioned earlier, this is unrealistic in our setting because in each period, the policy only observes a subset of r.v.s. So, we consider semi-bandit feedback, where the algorithm only observes the realizations of items that it probes. That is, at time $t$, the algorithm only observes $\left(X^{t}_{i}:i\in S(\sigma^{t},\mathbf{x}^{t})\right)$ where $S(\sigma^{t},\mathbf{x}^{t})$ is the set of items probed by policy $\sigma^{t}$ before it terminates.
1.2 Results and Techniques
@@
A stochastic problem $\mathcal{P}$with a minimization objective is up-monotone if for any instance $\mathcal{I}$ and probability distributions $\mathbf{{D}}$ and $\mathbf{{E}}$ where $\mathbf{{E}}$ stochastically dominates $\mathbf{{D}}$, we have $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{E}})\leq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$. Similarly, the problem is said to be down-monotone if for any instance $\mathcal{I}$ and probability distributions $\mathbf{{D}}$ and $\mathbf{{E}}$ where $\mathbf{{E}}$ stochastically dominates $\mathbf{{D}}$, we have $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{E}})\geq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$. The definitions for maximization problems are analogous.
@@
Suppose that a stochastic problem $\mathcal{P}$ has an $\alpha$-approximation algorithm, and it is either up-monotone or down-monotone. Then, there is a polynomial time semi-bandit learning algorithm for $\mathcal{P}$ (with unknown distributions) that has $\alpha$-regret $O(nkf_{max}\sqrt{T\log(kT)})$. Here, $n$ is the number of items, $k$ is the maximum support size of any distribution, $f_{max}$ is the maximal value of the objective function $f$ (over all realizations) and $T$ is the number of time periods.
We note that the dependence on the support-size $k$ can be replaced by the maximum number of “thresholds” used at any node of the stochastic policy (which corresponds to the number of branches out of the node). For most of our applications, we have approximate policies with $O(1)$ thresholds: so the regret bound is just $O(nf_{max}\sqrt{T\log T})$. See Theorem 2.9 for the formal statement, which also handles continuous r.v.s.
Consequently, we get polynomial time online learning algorithms for a number of stochastic optimization problems. We list a few of them next, highlighting the dependence on $n$ and $T$ (see §3 for further details and more applications).
  • Pandora’s box: $1$-regret $\widetilde{O}\left(n\sqrt{T\log T}\right)$
  • Prophet inequality: $1$-regret $\widetilde{O}\left(n\sqrt{T\log T}\right)$
  • Matroid Prophet inequality [KW19]: $\frac{1}{2}$-regret $\widetilde{O}\left(n\sqrt{T\log T}\right)$
  • Stochastic (maximization) knapsack [DGV08, Ma18]: $\frac{1}{2}$-regret $\widetilde{O}\left(n\sqrt{T\log T}\right)$
  • Stochastic covering knapsack [DHK16]: $3$-regret $\widetilde{O}\left(n\sqrt{T\log T}\right)$
The first two results improve over the regret bounds of $O\left(n^{3}\sqrt{T}\log(T)\right)$ and $O\left(n^{4.5}\sqrt{T}\log(T)\right)$ for prophet inequality and Pandora’s box from [GKSW24]. To the best of our knowledge, our results for all other applications are the first $\alpha$-regret bounds of $\sqrt{T}$. We note that the result of [GKSW24] for Pandora’s box and Prophet inequality also holds under the more restrictive bandit feedback model (only the objective value is observed). However, it is limited to these two problems, whereas our approach for semi-bandit feedback provides $\sqrt{T}$ regret for a much broader class of problems. Prior work (based on sample complexity) only gives $T^{2/3}$ regret for these problems [GHTZ21].
@@
The regret bound in Theorem 1.1 is nearly optimal in the following sense. There is an $\Omega(\sqrt{T})$ lower bound on the regret even in the much simpler setting of multi-armed-bandits [LR85]. Furthermore, if we move beyond our setting of independent-identically-distributed (i.i.d.) distributions across periods, to the “adversarial” setting with different distributions for each period, then there is a linear $\Omega(T)$ lower-bound on regret (even for prophet inequality) [GKSW24].
Technical Overview.
@@
At a high-level, our algorithm is based on the principle of optimism in the face of uncertainty which is well-studied in the multi-armed bandits literature: see e.g., the UCB algorithm [LS20]. Given observations from previous rounds, our algorithm first constructs a modified empirical distribution $\mathbf{{E}}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ which stochastically dominates the true distribution $\mathbf{{D}}$. It then executes the policy $\sigma$ given by the offline (approximation) algorithm under distribution $\mathbf{{E}}$.
@@
We now discuss how to bound the regret of the policy $\sigma$ under the true distribution $\mathbf{{D}}$. To keep things simple here, we assume that $\alpha=1$ and the problem is up-monotone. The difficulty in our setting is that the decision to probe/observe an item is random: it depends on the choice of policy $\sigma$ and the underlying distribution $\mathbf{D}$ of items. In contrast, the classical multi-armed bandits setting allows direct control to the algorithm on which item to probe. This gives algorithms such as UCB a direct control on the rate of exploration of individual items and allows them to bound the regret of individual items in terms of this rate. How do we bound the rate of exploration of different items in our setting when we do not have direct control on which items are probed? Moreover, how do we bound the regret contribution of individual items in terms of their rate of exploration? The key insight in our analysis is the following “stability” bound, which answers the above questions. Given product distributions $\mathbf{{D}}$ and $\mathbf{{E}}$, where $\mathbf{{E}}$ stochastically dominates $\mathbf{{D}}$, we show [^3] For policy $\sigma$ and distribution $\mathbf{D}$, $f(\sigma|\mathbf{{D}})$ is the expected objective of $\sigma$ when the r.v.s have distribution $\mathbf{D}$.
@@
where $\sigma^{*}$ (resp. $\sigma$) is an optimal policy under $\mathbf{{D}}$ (resp. $\mathbf{{E}}$), $Q_{i}(\sigma)$ is the probability that policy $\sigma$ probes item $i$, and $\epsilon_{i}$ is the total-variation distance between $\mathcal{E}_{i}$ and $\mathcal{D}_{i}$. The first inequality above follows from the monotone property. The second inequality requires analyzing the decision tree of $\sigma$ carefully and is a key technical contribution of this paper (see §2.2).

### v2

@@ Abstract
Stochastic optimization is a widely used approach for optimization under uncertainty, where uncertain input parameters are modeled by random variables. Exact or approximation algorithms have been obtained for several fundamental problems in this area. However, a significant limitation of this approach is that it requires full knowledge of the underlying probability distributions. Can we still get good (approximation) algorithms if these distributions are unknown, and the algorithm needs to learn them through repeated interactions? In this paper, we resolve this question for a large class of “monotone” stochastic problems, by providing a generic online learning algorithm with $\sqrt{T\log T}$ regret relative to the best approximation algorithm (under known distributions). Importantly, our online algorithm works in a semi-bandit setting, where in each period, the algorithm only observes samples from the random variables that were actually probed. Moreover, our result extends to settings with censored and binary feedback, where the policy only observes truncated or thresholded versions of the probed variables. Our framework applies to several fundamental problems such as prophet inequality, Pandora’s box, stochastic knapsack, single-resource revenue management and sequential posted pricing.
@@ 1 Introduction
Stochastic optimization problems have been a subject of intense investigation, as they offer a powerful lens through which we can handle uncertain inputs. In stochastic problems, uncertain input parameters are modeled by random variables, which are usually independent. Solutions (or policies) to a stochastic problem are sequential decision processes, where the next decision depends on all previously observed information. This approach has been applied to many domains: optimal stopping (Pandora’s box (Weitzman, 1979; Singla, 2018) and prophet inequality (Samuel-Cahn, 1984; Kleinberg and Weinberg, 2012; Feldman et al., 2021)), stochastic submodular optimization (Asadpour and Nazerzadeh, 2016; Golovin and Krause, 2017; Im et al., 2016), stochastic probing (Gupta and Nagarajan, 2013; Adamczyk et al., 2016; Gupta et al., 2017) and various other stochastic combinatorial optimization problems (e.g., knapsack (Deshpande et al., 2016; Dean et al., 2008) and matching (Chen et al., 2009; Bansal et al., 2012)). A fundamental assumption in all these results is that the underlying probability distributions are known to the algorithm. While the known-distributions assumption holds in some settings, it is not satisfied in many practical applications, e.g., in the absence of historical data. Our main goal in this paper is to relax this assumption and handle stochastic problems with unknown distributions. We start by providing some concrete examples.
@@
Series testing. There is a system with $n$ components, where each component $i$ is “working” independently with some probability $p_{i}$. All $n$ components must be working for the system to be functional. Moreover, it costs $c_{i}$ to test component $i$ and determine whether/not it is working. The goal is to test components sequentially to determine whether/not the system is functional, at the minimum expected cost. Clearly, testing will continue until some component is found to be not working, or we have tested all components and found them to be working. In the standard setting, where the probabilities $\{p_{i}\}_{i=1}^{n}$ are known upfront, the greedy policy that tests components in decreasing order of $\frac{1-p_{i}}{c_{i}}$ is optimal (Butterworth, 1972).
@@
Pandora’s box. There are $n$ items, where each item $i$ has an independent random value $\mathsf{X}_{i}$ with a known distribution $\mathcal{D}_{i}$. The realized value of $\mathsf{X}_{i}$ can be revealed by inspecting item $i$, which incurs cost $c_{i}$. The goal is to inspect a subset $S$ of the items (possibly adaptively) to maximize the expected utility $\mathbb{E}\left[\max_{i\in S}\mathsf{X}_{i}-\sum_{i\in S}c_{i}\right]$. We emphasize that a solution to this problem can be quite complex: the choice of the next item to inspect may depend on the realizations of all previous items. Nevertheless, there is an elegant optimal solution to this problem when the distributions $\{\mathcal{D}_{i}\}_{i=1}^{n}$ are known (Weitzman, 1979).
Single Resource Revenue Management. An airline has $C$ seats for sale across $n$ fare classes, where each seat sold in class $i$ has a fixed revenue $p_{i}$. The customer demand for each fare class $i$ is an independent random variable $\mathsf{X}_{i}\sim\mathcal{D}_{i}$. Customers preferring lower fares arrive earlier—a common assumption in practice. The airline allocates seats to different fare classes sequentially. If $m_{i}$ seats are offered to fare class $i$ then the associated revenue is $p_{i}\cdot\mathbb{E}[\min(X_{i},m_{i})]$. The number of seats offered to class $i$ may depend on the actual sales in prior classes. The objective is to maximize the expected total revenue. There is an exact dynamic programming algorithm for this problem (Gallego and Topaloglu, 2019).
Sequential Posted Pricing. There are $n$ buyers for an item, where each buyer $i$ has a random valuation $\mathsf{X}_{i}$ with known distribution $\mathcal{D}_{i}$. The seller wants to choose a take-it-or-leave-it price $p_{i}$ for each buyer $i$ so as to maximize the expected revenue when the buyers are offered these prices in some order. The first buyer $i$ whose value $\mathsf{X}_{i}$ exceeds their posted price $p_{i}$ buys the item, and the seller receives revenue $p_{i}$. There are two variants of this problem, depending on whether the buyer ordering is fixed or can be chosen by the algorithm. There is an exact dynamic program for the fixed-order setting and a $\left(1-\frac{1}{e}\right)$ approximation algorithm for the free-order setting (Yan, 2011).
@@
In each of the above problems, what if the underlying probabilities/distributions are unknown? Is it still possible to obtain good performance guarantees? In order to address this question formally, we utilize the framework of online learning. Here, the algorithm interacts with an unknown-but-fixed input distribution $\mathcal{D}$ over multiple time periods.In each period $t=1,\cdots T$, the online algorithm comes up with a solution/policy $\sigma^{t}$ to the stochastic problem and receives some feedback based on the performance of $\sigma^{t}$ on the (unknown) distribution $\mathcal{D}$.
@@
The type of feedback received is a crucial component in this learning-based framework. The simplest setting is full feedback, where the algorithm receives one sample from every random variable (r.v.). However, this is unrealistic for stochastic problems because the policy $\sigma^{t}$ in period $t$ only observes the r.v.s corresponding to some subset of the items. For example, in series testing, $\sigma^{t}$ tests the $n$ components in some order until the first non-working component is found: if the first k1k-1 components are working and the $k^{th}$ component is not working then we would only observe the outcomes/samples of the first $k$ r.v.s (and the remaining nkn-k r.v.s are unobserved). In this paper we consider the more natural setting of semi-bandit feedback, where in each period $t$, the algorithm only receives samples from the r.v.s that the policy $\sigma^{t}$ actually observed. We further study more restricted feedback models— censored and binary feedback—that naturally arise in applications like revenue management and posted pricing, respectively.
@@
Our goal is to minimize the expected regret of the online algorithm, which is the difference between the total $T$-period objective of our algorithm and the optimum (which knows the distribution $\mathcal{D}$). Obtaining $o(T)$ regret dependence with respect to $T$ means that our online algorithm, which doesn’t know $\mathcal{D}$, asymptotically approaches the optimum. An additional difficulty arises from the fact that many problems that we consider are NP-hard (e.g., knapsack and submodular optimization). In such cases, we cannot hope to approach the optimum (via a polynomial algorithm). Here, we use the notion of $\alpha$-regret, where the online algorithm approaches the $\alpha$ approximately optimal value. Our main result is a general method for transforming any offline $\alpha$-approximation algorithm (with known distributions) to an online learning algorithm with $\alpha$-regret of $\widetilde{O}(\sqrt{T})$. It is well-known that the $\sqrt{T}$ regret bound cannot be improved, even in very special cases. Our method works for a wide range of stochastic problems that satisfy a natural monotonicity condition. This includes all the above examples as well as several other fundamental stochastic optimization problems.
@@ 1.1 Problem Setup
Stochastic Optimization
Consider a stochastic problem $\mathcal{P}$ where the input consists of $n$ items with an independent real-valued r.v. $\mathsf{X}_{i}$ associated with each item $i\in[n]=\{1,2,\cdots n\}$. There is a known probability distribution $\mathcal{D}_{i}$ for $\mathsf{X}_{i}$; that is, $\mathsf{X}_{i}\sim\mathcal{D}_{i}$. We will primarily consider the setting where each $\mathcal{D}_{i}$ has finite support, and denote the set of outcomes of all r.v.s by $O$. (We also consider continuous distributions in §3.) In order to determine the realization of any r.v. $\mathsf{X}_{i}$, we need to select or probe item $i$.
@@
The cost of policy $\sigma$ depends on the realization $\mathbf{x}$, and we use $f(\sigma,\mathbf{x})\geq 0$ to denote the cost of policy $\sigma$ under realization $\mathbf{x}$. Specifically, $f(\sigma,\mathbf{x})$ depends only on the sequence of probed items $S(\sigma,\mathbf{x})$ and their realizations. We assume that cost is accrued only at terminal/leaf nodes: this is without loss of generality as every policy execution ends at a leaf node. [^1] In some problems, it is natural to accrue cost at each state that the policy goes through. We can easily convert such a cost structure to our setting by placing all the cost accrued from a policy execution at its terminal state. We make no further assumptions about the cost function $f$ beyond boundedness: it does not have to be linear, submodular etc. Our goal is to find a policy $\sigma$ that minimizes the expected cost $f(\sigma):=\mathbb{E}_{\mathbf{x}}[f(\sigma,\mathbf{x})]$. We also allow for constraints in problem $\mathcal{P}$, which must be satisfied under all realizations (i.e., with probability $1$). Let $\mathcal{C}$ denote the set of feasible policies (that satisfy the constraints in $\mathcal{P}$). Sometimes, we work with randomized policies, where each node in the decision tree corresponds to a probability distribution over items (rather than a single item). Randomized policies are not any stronger than deterministic ones: there is always an optimal deterministic policy for this class of problems. The stochastic problem $\mathcal{P}$ is as follows (maximization problems are defined similarly).
@@
Observe that the number of nodes in policy $\sigma$ may be exponential. So, we are interested in “efficient” policies that can be implemented in polynomial-time for any realization $\mathbf{x}$. For some stochastic problems like series systems and Pandora’s box, efficient optimal policies are known. On the other hand, there are many problems like stochastic knapsack, matching and set-cover, where optimal policies may not be efficient: in these cases, we will focus on (efficient) approximately optimal policies. We use $\sigma^{*}$ to denote an optimal policy, and denote its expected cost by $\mathtt{OPT}$. An $\alpha$-approximation algorithm for a stochastic problem $\mathcal{P}$ takes as input the probability distributions $\{\mathcal{D}_{i}\}_{i=1}^{n}$ (along with any objective/constraint parameters) and returns a policy that has expected cost at most $\alpha\cdot\mathtt{OPT}$. We use the convention that for minimization problems, the approximation ratio $\alpha\geq 1$, whereas for maximization problems $\alpha\leq 1$.
The Online Semi-Bandit Setting
In this setting, the distributions $\mathcal{D}_{1},\ldots,\mathcal{D}_{n}$ are unknown (other parameters such as the cost function and constraints are known). Moreover, we have to repeatedly solve the stochastic problem $\mathcal{P}$ over $T$ time periods. We assume throughout the paper that $T\geq n$. The distributions $\{\mathcal{D}_{i}\}_{i=1}^{n}$ remain the same across all $T$ periods. The goal is to simultaneously learn the distributions and converge to a good policy over time. At time $t\leq T$, we use all prior observations to present policy $\sigma^{t}$. The policies $\sigma^{t}$ may be different for each time $t\in[T]$. We measure our learning algorithm in terms of expected total regret. For minimization problems, we define it as follows.
@@
$R(T)$ $=\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\mathbb{E}_{\mathbf{x}^{t}}\left[f(\sigma^{t},\mathbf{x}^{t})\right]-T\cdot\mathtt{OPT}.$ (2)
@@
We consider the setting of semi-bandit feedback, where the algorithm only observes the realizations of items that it probes. That is, at time $t$, the algorithm only observes $\left(x^{t}_{i}:i\in S(\sigma^{t},\mathbf{x}^{t})\right)$ where $S(\sigma^{t},\mathbf{x}^{t})$ is the set of items probed by policy $\sigma^{t}$ before it terminates.
Censored/Binary Feedback
We also extend our results to more restricted feedback settings, where the policy may not be able to directly observe $\mathsf{X}_{i}$ for a probed item $i$. In particular, we consider two such extensions. In censored feedback, the policy only observes the truncated realization $\min(x_{i},\tau)$ where $\tau$ is some threshold. This is useful in revenue management applications with random demand. In binary feedback, the policy only observes whether or not the realization $x_{i}$ is more than the threshold $\tau$. This is useful in posted pricing and mechanism design.
1.2 Main Result and Techniques
@@
A stochastic problem $\mathcal{P}$ is up-monotone if for any instance $\mathcal{I}$ and probability distributions $\mathbf{{D}}$ and $\mathbf{{E}}$ where $\mathbf{{E}}$ stochastically dominates $\mathbf{{D}}$, we have $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{E}})\leq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$. Similarly, the problem is said to be down-monotone if for any instance $\mathcal{I}$ and probability distributions $\mathbf{{D}}$ and $\mathbf{{E}}$ where $\mathbf{{E}}$ stochastically dominates $\mathbf{{D}}$, we have $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{E}})\geq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$.
@@
Suppose that a stochastic problem $\mathcal{P}$ has an $\alpha$-approximation algorithm, and it is either up-monotone or down-monotone. Then, there is a polynomial time semi-bandit learning algorithm for $\mathcal{P}$ (with unknown distributions) that has $\alpha$-regret $O(nkf_{max}\sqrt{T\log(nT)})$. Here, $n$ is the number of items, $k$ is the maximum support size of any distribution, $f_{max}$ is the maximal value of the objective function $f$ (over all realizations) and $T$ is the number of time periods.
If we do not care about polynomial running time then we can just set $\alpha=1$ (by computing the optimal policy for $\mathcal{P}$) and obtain $1$-regret $O(nkf_{max}\sqrt{T\log(nT)})$ for any monotone stochastic problem. The only general result of this type known prior to our work is a $T^{2/3}$ regret bound for these problems using sample complexity bounds of Guo et al. (2021).
@@
The regret bound in Theorem 1.1 is nearly optimal in the following sense. There is an $\Omega(\sqrt{T})$ lower bound on the regret even in a very special case: see Appendix B. Furthermore, if we move beyond our setting of independent-identically-distributed (i.i.d.) distributions across periods, to the “adversarial” setting with different distributions for each period, then there is a linear $\Omega(T)$ lower-bound on regret for prophet inequality (Gatmiry et al., 2024), which shows that sublinear regret is not possible for our class of problems in the adversarial case.
@@
Technical Overview. At a high-level, our algorithm is based on the principle of optimism in the face of uncertainty which is well-studied in the multi-armed bandits literature: see e.g., the UCB algorithm (Auer et al., 2002). Given observations from previous rounds, our algorithm first constructs a modified empirical distribution $\mathbf{{E}}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ which stochastically dominates the true distribution $\mathbf{{D}}$. It then executes policy $\sigma$ given by the offline (approximation) algorithm under distribution $\mathbf{{E}}$.
@@
We now discuss how to bound the regret of the policy $\sigma$ under the true distribution $\mathbf{{D}}$. To keep things simple here, we assume that $\alpha=1$. We also assume that the problem is up-monotone and has a minimization objective (our analysis for down-monotonicity and/or maximization objective is identical). The difficulty in our setting is that the decision to probe/observe an item is random: it depends on the choice of policy $\sigma$ and the underlying distribution $\mathbf{D}$ of items. In contrast, the classical multi-armed bandits setting allows direct control to the algorithm on which item to probe: so the UCB algorithm can control the rate of exploration of individual items and bound the regret of individual items in terms of this rate. How do we bound the rate of exploration of different items when we do not have direct control on which items are probed? Moreover, how do we bound the regret contribution of individual items in terms of their rate of exploration?
The key insight in our analysis is the following “stability” bound, which answers the above questions. Given product distributions $\mathbf{{D}}$ and $\mathbf{{E}}$, where $\mathbf{{E}}$ stochastically dominates $\mathbf{{D}}$, we show [^2] For policy $\sigma$ and distribution $\mathbf{D}$, $f(\sigma|\mathbf{{D}})$ is the expected objective of $\sigma$ when the r.v.s have distribution $\mathbf{D}$.
@@
where $\sigma^{*}$ (resp. $\sigma$) is an optimal policy under $\mathbf{{D}}$ (resp. $\mathbf{{E}}$), $Q_{i}(\sigma)$ is the probability that policy $\sigma$ probes item $i$, and $\epsilon_{i}$ is the total-variation distance between $\mathcal{E}_{i}$ and $\mathcal{D}_{i}$. The first inequality above follows from the monotonicity property. The second inequality requires analyzing the decision tree of $\sigma$ carefully and is a key technical contribution of this paper (see §2.1).